# No Images, No Problem: Retaining Knowledge in Continual VQA with Questions-Only Memory

### v1

@@ 3.4 Attention Consistency Distillation
Continual VQA demands a careful balance between stability and plasticity, particularly as models must adapt to new tasks without forgetting. While pseudo-labeling offers a straightforward way to preserve prior knowledge, it provides limited stability because it only influences the network’s outputs, leaving its internal representations vulnerable to drift. Feature distillation across multiple layers [29, 77, 11, 54] provides stronger regularisation by guiding internal representations but often compromises plasticity. These methods impose rigid constraints across all layers, which is particularly problematic in our setting, where new image-question pairs $(x^{t},q^{m})$ introduced during learning the current task were not seen in previous tasks. Although these pairs are meaningful and aligned, the model has no prior experience processing them, making strict layer-wise regularisation overly restrictive. Our VQA model [85, 52] processes image features and language tokens as a unified sequence in a transformer, where self-attention naturally captures both intra-modal (language-to-language and image-to-image) and inter-modal (language-to-image) interactions. This implicit cross-attention capability enables the model to compute similarity scores between language and image tokens without requiring cross-attention layers. However, as the model is sequentially fine-tuned on new tasks, it runs the risk of suffering from self-attention drift—a gradual shift in focus away from relevant visual regions associated with previous tasks [65, 21]. This drift occurs because pseudo-labeling alone cannot prevent changes in the model’s internal representations.
To address this, we introduce an intermediate approach: attention consistency distillation. Our approach aligns the self-attention weights of the current model $\phi^{t}$ with those of the previous task model $\phi^{t-1}$, preserving both intra-modal and inter-modal relationships. This alignment acts as a lightweight but effective regularisation that maintains stability within the self-attention patterns. By distilling only the attention maps, we constrain the model to maintain its focus on relevant visual regions for answering questions without imposing strong regularisation on all layers.
Let us sample an image-question pair $(x^{t},q^{m})$ with ${}{x^{t}\sim\mathcal{T}^{t}},~{}{q^{m}\sim\mathcal{M}}$: we denote $A_{k}^{t}$ and $A_{k}^{t-1}$ as the corresponding self-attention maps in an attention head $k$ for the current and previous models, respectively, computed over the unified token sequence of both language and image features [85, 52]. These maps are generated by applying softmax to the scaled query-key product, yielding normalised distributions that represent the focus between intra-modal and inter-modal interactions. To align these distributions, we define the self-attention distillation loss $\mathcal{L}_{\text{Att}}$ using cross-entropy, encouraging the current model’s attention maps $A_{k}^{t}$ to resemble those from the previous task $A_{k}^{t-1}$:
$\begin{split}\mathcal{L}_{\text{Att}}=\mathbb{E}_{x^{t}\sim\mathcal{T}^{t}}\mathbb{E}_{q^{m}\sim\mathcal{M}}&\mathbb{E}_{k\sim\mathcal{K}_{\phi}}\\ &\mathcal{L}_{\text{CE}}\left[A_{k}^{t}(x^{t},q^{m}),A_{k}^{t-1}(x^{t},q^{m})\right],\end{split}$ (5)
where $\mathcal{K}_{\phi}$ denotes the set of all the attention heads across all the layers of $\phi$. Our approach employs cross-entropy loss over normalized attention maps, in contrast to prior works [11, 54] that rely on L1 or ReLU+L1 losses. Cross-entropy enforces alignment between attention distributions, preserving both individual token importance and structured inter-modal relationships essential for VQACL. In contrast, L1-based methods treat attention scores as independent values, imposing rigid constraints that limit adaptability to new tasks. A key advantage of cross-entropy is its consistency across all loss terms in QUAD ($\mathcal{L}_{\text{CE}},\mathcal{L}_{\text{PL}},\mathcal{L}_{\text{Att}}$), ensuring coherent optimization. Empirically, this design mitigates catastrophic forgetting.
@@ 4.1 Experimental setup
Implementation Details. To ensure a fair comparison, we adopt the protocol of [85] for both feature extraction and training across datasets. For the visual embeddings, we use a Faster R-CNN [58] trained on the Visual Genome dataset [34] to obtain 36 region-based features per image in the VQAv2 dataset. For videos in the NExT-QA dataset, we extract clip-level motion features using an inflated 3D ResNeXt-101 [25], setting $n=16$ regions per clip. These visual features are adapted through a two-layer MLP with GELU activation, preparing them as inputs to the transformer model. Our transformer backbone, based on T5 [55], consists of 12 blocks for both encoder and decoder modules, each containing 12 attention heads. The embedding dimension $d$ is tailored to the task-specific requirements. Training is conducted for 3 epochs per task, with a batch size of 80. We utilize the Adam optimizer [31] with an initial learning rate of $10^{-4}$. $\lambda$ is set to $0.5$ in all experiments. All implementations are based on PyTorch [53]. Unlike VQACL [85], QUAD does not create or store any prototypes for questions or visual objects.
@@
Evaluation Metrics. We utilise two established metrics for continual learning [85, 4, 45]: final average performance ($AP$), and average forgetting ($Forget$). The $AP$ metric reflects the model’s overall performance across all learnt tasks, highlighting its ability to consistently acquire new tasks. Let $a_{i,j}$ represent the performance of the model on task $\mathcal{T}^{i}$ after it has completed learning this task $\mathcal{T}^{j}$. Then, $AP$ is calculated as: $AP\!=\!\frac{1}{T}\sum_{t=1}^{T}a_{t,T}.$ Additionally, the $Forget$ metric quantifies the performance loss in previous tasks as new tasks are learnt and is computed as: $Forget\!=\!\frac{1}{T-1}\sum_{t=1}^{T-1}\max_{z\in\{t,\ldots,T-1\}}(a_{t,z}-a_{t,T}).$ To ensure a fair comparison, we calculate $a_{i,j}$ in the NExT-QA dataset as per the approach in [85, 75], utilising Wu-Palmer Similarity (WUPS) [74] to assess the quality of the generated answers. For the VQAv2 dataset, following the method outlined in [85], we use the percentage of correctly answered questions as the value for $a_{i,j}$.
@@
Baselines. To rigorously evaluate the performance of our method, we conduct a comprehensive analysis encompassing five established and state-of-the-art continual learning methods. These methods include two regularisation-based approaches: Elastic Weight Consolidation (EWC) [33] and Memory Aware Synapses (MAS) [1], as well as three rehearsal-based approaches: Experience Replay (ER) [7], Dark Experience Replay (DER) [6], Virtual Sample (VS) [67], and VQACL [85] (see details in Appendix). To contextualise the performance of these methods, we establish two performance boundaries: a lower bound (Vanilla) that performs simple gradient updates without any forgetting mitigation strategies, and an upper bound (Joint) that trains on all tasks simultaneously, providing an upper bound performance metric. Following the evaluation protocol established in VQACL [85], we employ two key evaluation strategies: standard testing, which evaluates the models’ performance on previously encountered task types, assessing their ability to retain learnt knowledge; and novel composition testing, which challenges the models with previously unseen combinations of visual and linguistic elements, probing their capacity for compositional generalization. This dual-pronged evaluation strategy, adopted from the original VQACL work, provides a nuanced understanding of each method’s efficacy in balancing knowledge retention and adaptive learning in the VQACL.
@@ 5 Ablation Study and Analysis
Figure 3: Plasticity/Stability analysis on VQAv2. Each matrix shows the performance of a model trained on tasks (rows) and evaluated on tasks sequentially (columns). The diagonal (highlighted in orange) represents in-domain performance, while off-diagonal elements indicate cross-domain generalization. Higher values (darker colors) suggest better retention and plasticity. The progression from ‘Sequential finetuning’, to $\mathcal{L}_{PL}$ and then to our full method, QUAD, highlights the improvement in retaining knowledge across sequential tasks.
@@
Distillation Components. Tab. 3 shows the impact of PL distillation and attention distillation in QUAD. We evaluate three setups: (1) PL distillation alone, (2) attention distillation alone, and (3) the combination of both.
Using PL distillation alone achieves moderate performance, with AP of 30.72% on VQAv2 and 29.04% on NExT-QA, indicating effective feature alignment across tasks. In contrast, attention distillation alone yields lower AP scores (13.34% on VQAv2 and 13.24% on NExT-QA) with high forgetting scores (32.08% on VQAv2 and 24.56% on NExT-QA), suggesting limited task adaptation. Combining PL and attention distillation achieves the best results, with AP scores of 39.25% on VQAv2 and 31.70% on NExT-QA, and the lowest forgetting rates.
@@
Attention Distillation. Tab. 4 demonstrates the effectiveness of our proposed $\mathcal{L}_{\text{Att}}$ within QUAD. Unlike prior methods such as Attn-dist (L1) and Asym-Attn [54], which impose L1 or ReLU+L1 losses on raw attention scores, our approach applies cross-entropy over normalized attention maps. QUAD outperforms previous methods, achieving an AP of 39.25% and Forgetting of 4.91% on VQAv2, and an AP of 31.70% with Forgetting of 2.91% on NExT-QA. These results highlight the importance of preserving distributional consistency in attention maps and avoiding unstructured alignment.
Table 3: Ablation study of QUAD components on VQAv2 and NExT-QA.
$\mathcal{L}_{\text{PL}}$ $\mathcal{L}_{\text{Att}}$ Memory VQAv2 NExT-QA
@@
$\mathcal{L}_{\text{PL}}$ + Attn-dist (L1) 34.56 7.91 30.14 5.78
$\mathcal{L}_{\text{PL}}$ + Asym-Attn [54] 38.15 5.57 31.18 4.13
@@
Analysing Plasticity/Stability. Fig. 3 illustrates the impact of different strategies on plasticity and forgetting in the VQAv2 dataset. The sequential fine-tuning matrix (left) reveals the impact of learning without rehearsal, showing uniformly low off-diagonal values—indicating the model’s failure to answer questions from previous tasks. This phenomenon, which we term the out-of-answer-set problem, occurs when the model overfits to the current task’s answer space, preventing it from correctly responding to questions from earlier tasks. The $\mathcal{L}_{\text{PL}}$ matrix (center) mitigates forgetting through pseudo-label distillation, significantly improving cross-domain retention in tasks such as “commonsense” and “count.” However, it remains less effective in complex reasoning tasks like “causal” and “subcategory.” Our full method, QUAD (right), achieves the best overall performance by integrating pseudo-label and attention consistency distillation, striking a balance between stability and plasticity. QUAD consistently maintains high values both on and off the diagonal, demonstrating superior retention of prior knowledge and adaptability to new tasks. Attention consistency can be assessed by comparing the right and middle subplots. For instance, after training with QUAD, the model retains more knowledge on the ‘judge’ task (62.6%) compared to pseudo-labeling alone (35.0%).
@@
Sensitivity to Memory Size. Fig. 4 shows a sensitivity analysis of CL methods with respect to memory size. QUAD consistently outperforms baseline methods (ER, DER, VS, VQACL) across all memory sizes, demonstrating the effectiveness of our distillation strategy. On VQA-v2, QUAD exhibits strong scalability, with AP increasing steadily from 1000 to 5000 samples. In constrained storage scenarios, QUAD maintains competitive performance by leveraging question-only distillation. On NExT-QA, QUAD outperforms other methods, though with smaller gains than on VQAv2 due to the increased task complexity and reliance on visual information.
@@
Effectiveness of Object-Matched Question Selection. Fig. 5 demonstrates the benefits of object-matched question selection over random selection. We analyze its impact across memory sizes on VQAv2 and NExT-QA, where it consistently improves AP as memory grows. Aligning selected questions with object categories enhances knowledge transfer, leading to better adaptation to new tasks.
@@ 6 Conclusion
In this work, we introduced QUestion-only replay with Attention Distillation (QUAD), a novel distillation-based approach for continual VQA. Unlike conventional methods that store both images and questions, QUAD addresses storage and privacy concerns by retaining only past task questions. This design enables effective regularization without storing sensitive visual data, making it highly practical for privacy-conscious applications.
Comprehensive evaluations on VQAv2 and NExT-QA show that QUAD consistently outperforms both memory-free and memory-rehearsal methods, achieving state-of-the-art generalization and knowledge retention. These results validate question-only replay as an efficient and scalable solution for continual VQA. Future work could explore dynamically generated questions to eliminate memory storage entirely, further enhancing adaptability in real-world multimodal scenarios.
Acknowledgements. This paper has been supported by the French National Research Agency (ANR) in the framework of its JCJC, and was funded by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS). Furthermore, this research was partially funded by Hi!PARIS Center on Data Analytics and Artificial Intelligence. This work was granted access to the HPC resources of IDRIS under the allocation AD011013860 made by GENCI.
@@
In this supplementary material, we provide additional details on the experimental results presented in the main paper, along with further empirical evaluations and discussions.
Section 7 presents an ethics statement regarding our work. Section 8 discusses the limitations of our proposed method, QUAD. Section 9 reports the computational and memory footprint analysis of QUAD. Section 10 and Section 11 describe the datasets, task sequences, and evaluation structure. Section 12 provides analysis on plasticity-stability for NExT-QA dataset. Lastly, Section 13 explains the baselines used for comparison.

### v2

@@ 3.4 Attention Consistency Distillation
While question-only replay using pseudo-labeling ensures output consistency, it fails to constrain internal representations, causing self-attention drift and disrupting alignment with previously learned tasks [24, 79, 63]. In contrast, feature distillation across multiple layers [34, 87, 14, 63] offers stronger regularization by aligning internal representations; however, it often restricts plasticity due to its rigid constraints. These methods enforce layer-wise alignment, which becomes problematic when encountering new image-question pairs $(x^{t},q^{m})$ that were not present in previous training. Despite their semantic relevance, the model lacks prior exposure to these pairs, making strict layer-wise regularization overly restrictive. Our VQA model [95, 61] encodes image and text tokens within a unified transformer sequence, allowing self-attention to dynamically capture intra-modal (text-to-text, image-to-image) and inter-modal (text-to-image) dependencies. However, sequential fine-tuning gradually shifts self-attention [24, 79], causing the model to focus on different visual regions than those relevant to past tasks, ultimately degrading performance on prior knowledge [75, 24]. This drift arises because pseudo-labeling alone fails to stabilize internal representations, necessitating explicit regularization.
To mitigate this, we introduce attention consistency distillation (ACD), which aligns the self-attention maps of the current model $\phi^{t}$ and its predecessor $\phi^{t-1}$. By preserving intra-modal and inter-modal relationships, this regularization stabilizes attention patterns while maintaining flexibility. Distilling only attention maps guides the model to attend to task-relevant visual regions for question answering while preserving adaptability to novel input patterns.
For an image-question pair $(x^{t},q^{m})$ with $x^{t}\sim\mathcal{T}^{t},q^{m}\sim\mathcal{M}$: we denote $A_{k}^{t}$ and $A_{k}^{t-1}$ as the corresponding self-attention maps in an attention head $k$ for the current and previous model. Thus, we define the self-attention consistency loss $\mathcal{L}_{\text{ACD}}$ via cross-entropy, ensuring that attention distributions remain aligned across tasks:
$\begin{split}\mathcal{L}_{\text{ACD}}=\mathbb{E}_{x^{t}\sim\mathcal{T}^{t}}\mathbb{E}_{q^{m}\sim\mathcal{M}}&\mathbb{E}_{k\sim\mathcal{K}_{\phi}}\\ &\mathcal{L}_{\text{CE}}\left[A_{k}^{t}(x^{t},q^{m}),A_{k}^{t-1}(x^{t},q^{m})\right],\end{split}$ (5)
where $\mathcal{K}_{\phi}$ represents all attention heads across layers of $\phi$.
Unlike prior L1-based approaches [14, 63], which impose uniform penalties and operate directly on the raw query-key products, ACD leverages cross-entropy on normalized attention maps (after softmax operation) to preserve the probabilistic structure of attention distributions. This gives greater emphasis to highly attended regions while low-attended areas remain flexible, preventing over-penalization of less critical regions. Asymmetric regularization (e.g., ReLU+L1)  [63] partially addresses this issue by prioritizing reductions in attended regions but lacks a probabilistic alignment mechanism, making it less effective in maintaining structured dependencies in multimodal learning. In contrast, ACD offers an importance-weighted alignment that preserves essential cross-modal associations without hindering adaptation. The effectiveness of ACD is empirically demonstrated in Sec. 4, and further discussed in the Appendix.
@@ 4.1 Experimental setup
Implementation Details. To ensure a fair comparison, we adopt the protocol of [95] for both feature extraction and training across datasets. For the visual embeddings, we use a Faster R-CNN [67] trained on the Visual Genome dataset [40] extracting 36 region-based object features per image in the VQAv2 dataset. For videos in the NExT-QA dataset, we extract clip-level motion features using an inflated 3D ResNeXt-101 [29], setting $n=16$ regions per clip. A two-layer MLP with GELU activation adapts these features for input into the transformer backbone. Our transformer backbone, based on T5 [64], consists of 12 blocks for both encoder and decoder modules, each containing 12 attention heads. The embedding dimension $d$ is tailored to the task-specific requirements. Training is conducted for 3 epochs per task, with a batch size of 80. We utilize the Adam optimizer [36] with an initial learning rate of $10^{-4}$. $\lambda$ is set to 0.50.50.5 in all experiments. All implementations are based on PyTorch [62]. Importantly, unlike [95], QUAD does not store any visual or question prototypes.
@@
Evaluation Metrics. We utilise two established metrics for continual learning [95, 6, 53]: final average performance ($AP$), and average forgetting ($Forget$). The $AP$ metric reflects the model’s overall performance across all learnt tasks, highlighting its ability to consistently acquire new tasks. Let $a_{i,j}$ represent the performance of the model on task $\mathcal{T}^{i}$ after it has completed learning this task $\mathcal{T}^{j}$. Then, $AP$ is calculated as: $AP\!=\!\frac{1}{T}\sum_{t=1}^{T}a_{t,T}.$ Additionally, the $Forget$ metric serves as a proxy for knowledge retention, quantifying performance degradation on prior tasks as new tasks are learned. It is computed as: $Forget\!=\!\frac{1}{T-1}\sum_{t=1}^{T-1}\max_{z\in\{t,\ldots,T-1\}}(a_{t,z}-a_{t,T}).$ To ensure a fair comparison, For NExT-QA, following [95, 85], we compute $a_{i,j}$ using Wu-Palmer Similarity (WUPS) to evaluate answer quality. For the VQAv2 dataset, as described in [95], we use the percentage of correctly answered questions as the value for $a_{i,j}$.
@@
Baselines. We benchmark QUAD against five established continual learning methods spanning regularization and rehearsal-based strategies. This includes two regularization-based methods: Elastic Weight Consolidation (EWC) [38] and Memory Aware Synapses (MAS) [2], as well as three rehearsal-based approaches: Experience Replay (ER) [9], Dark Experience Replay (DER) [8], Virtual Sample (VS) [77], and VQACL [95] (see details in Appendix). We additionally report lower and upper performance bounds to frame the results: a lower bound (Vanilla) using naive finetuning without forgetting mitigation, and an upper bound (Joint) that trains on all tasks simultaneously. Following the evaluation protocol in VQACL [95], we employ two key evaluation strategies: standard testing, which evaluates the models’ performance on previously encountered task types, assessing their ability to retain learnt knowledge, and novel composition testing, which challenges the models with previously unseen combinations of visual and linguistic elements, probing their capacity for compositional generalization. This twofold evaluation reveals how well each method balances retention and adaptation (details in Appendix). Additional experiments using pretrained models are included in the Supp Mat.
@@ 5 Ablation Study and Analysis
Figure 4: Plasticity/Stability analysis on VQAv2. Each matrix shows the performance of a model trained on tasks (rows) and evaluated on tasks sequentially (columns). The diagonal (highlighted in orange) represents in-domain performance, while off-diagonal elements indicate cross-domain generalization. Higher values (darker colors) suggest better retention and plasticity. The progression from ‘Sequential finetuning’, to $\mathcal{L}_{QR}$ and then to our full method, QUAD, highlights the improvement in retaining knowledge across sequential tasks.
@@
QUAD Components. Tab. 3 shows the impact of questions-only replay and attention distillation in QUAD. We evaluate: (1) Question-only replay alone, (2) attention distillation alone, and (3) the combination of both.
Using question-only replay mechanism achieves moderate performance, with AP of 30.72% on VQAv2 and 29.04% on NExT-QA, indicating effective knowledge preservation across tasks. In contrast, attention distillation alone yields lower AP scores (13.34% on VQAv2 and 13.24% on NExT-QA) with high forgetting scores (32.08% on VQAv2 and 24.56% on NExT-QA), suggesting limited task adaptation. Combining $\mathcal{L}_{\text{QR}}$ and $\mathcal{L}_{\text{ACD}}$ achieves the best results, with AP scores of 39.25% on VQAv2 and 31.70% on NExT-QA, and the lowest forgetting rates.
@@
Attention Distillation. Tab. 4 demonstrates the effectiveness of our proposed $\mathcal{L}_{\text{ACD}}$ within QUAD. Unlike prior methods such as Attn-dist (L1) and Asym-Attn [63], which impose L1 or ReLU+L1 losses on raw attention scores, our approach applies cross-entropy over normalized attention maps. QUAD outperforms previous methods, achieving an AP of 39.25% and Forgetting of 4.91% on VQAv2, and an AP of 31.70% with Forgetting of 2.91% on NExT-QA. These results highlight the benefit of distributional consistency in attention maps over unstructured alignment.
Table 3: Ablation study of QUAD components.
$\mathcal{L}_{\text{QR}}$ $\mathcal{L}_{\text{ACD}}$ Memory VQAv2 NExT-QA
@@
$\mathcal{L}_{\text{QR}}$ + Attn-dist (L1) 34.56 7.91 30.14 5.78
$\mathcal{L}_{\text{QR}}$ + Asym-Attn [63] 38.15 5.57 31.18 4.13
@@
Analysing Plasticity/Stability. Fig. 4 compares how different methods manage forgetting and adaptation on the VQAv2 dataset. The sequential fine-tuning baseline (left) exhibits uniformly low off-diagonal scores, signaling a complete failure to retain knowledge from earlier tasks. This severe forgetting stems from overfitting to the current task’s answer space, a phenomenon we call the out-of-answer-set problem, it occurs when the model overfits to the current task’s answer space, preventing it from correctly responding to questions from earlier tasks. By contrast, question-only replay $\mathcal{L}_{\text{QR}}$ (center) noticeably improves retention through our question-only replay mechanism, especially in tasks like “commonsense” and “count”. However, it remains less effective in complex reasoning tasks like “causal” and “subcategory”.
Our full method, QUAD (right), achieves the best balance: it sustains high diagonal accuracy while boosting off-diagonal retention. reflecting both adaptability and long-term memory. For example, QUAD preserves 62.6% accuracy on the ‘judge’ task, compared to only 35.0% with replay alone—highlighting the value of attention consistency distillation. Notably, tasks like ‘type’ that rely heavily on visual semantics remain challenging under the question-only setting, reaffirming the need for richer visual grounding in some categories. Nevertheless, QUAD excels in conceptually driven tasks such as “commonsense”, showing its effectiveness even with reduced supervision.
@@
Sensitivity to Memory Size. Fig. 5 shows how different continual learning methods respond to varying memory budgets. Across all memory sizes, QUAD consistently outperforms baselines (ER, DER, VS, VQACL), demonstrating the effectiveness of our distillation strategy. On VQAv2, QUAD exhibits robust scalability, with AP increasing steadily from 1K to 5K samples. In constrained storage scenarios, QUAD maintains competitive performance by leveraging question-only distillation. On the more visually complex NExT-QA benchmark, QUAD still leads, though with narrower margins. This reflects the greater challenge of retaining visual-semantic alignment using question-only signals.
@@
Effectiveness of Object-Matched Question Selection. Fig. 6 highlights the advantage of object-matched question selection compared to random pairing. We analyze its impact across memory sizes on VQAv2 and NExT-QA, where it consistently improves AP as memory grows. This targeted selection ensures semantic relevance between the question and visual context, which helps reinforce meaningful cross-task associations. As a result, the model adapts effectively to new tasks while retaining prior knowledge.
@@ 6 Conclusion
In this work, we introduced QUestion-only replay with Attention Distillation (QUAD), a novel questions-only replay framework for continual VQA. Unlike conventional methods that store both images and questions, QUAD addresses storage and privacy concerns by retaining only past task questions. This design enables effective regularization without storing sensitive visual data, making it highly practical for privacy-conscious applications. Comprehensive evaluations on VQAv2 and NExT-QA show that QUAD consistently outperforms both memory-free and memory-rehearsal methods, achieving state-of-the-art performance. Surprisingly, our method, without any image exemplars, outperforms previous methods, which do require image storage.
Acknowledgements. This paper is supported by the French National Research Agency (ANR) in the framework of the JCJC project “BANERA under Grant ANR-24-CE23-4369, and was funded by the European Union’s Horizon Europe research and innovation program under grant agreement No. 101120237 (ELIAS). It was also partially funded by Hi!PARIS Center on Data Analytics and Artificial Intelligence and was granted access to the HPC resources of IDRIS under the allocation AD011013860 made by GENCI.
@@
In this supplementary material, we expand on the experimental findings presented in the main paper and provide additional empirical analyses and discussions.
Section 7 outlines the ethical considerations of our work, while Section 8 discusses the limitations of our proposed method, QUAD. Sections 9 and 10 delve into the benefits of our attention consistency distillation (ACD) approach compared to conventional L1-based attention regularization. In Section 11, we present a detailed analysis of the computational and memory footprint of QUAD. Section 12 examines the impact of the balancing hyperparameter $\lambda$. Additional results using pretrained vision-language models (BLIP-2 and LLaVA-7B) are reported in Section 13.
Section 14 provides further insights into the out-of-answer-set problem encountered in continual VQA. Sections 15 and 16 describe the datasets used, task orderings, and evaluation protocol. Section 17 offers an extended analysis of the plasticity-stability trade-off on the NExT-QA dataset. Finally, Section 18 details the baseline methods used for comparison throughout our study.